OpenStudy (anonymous):

Which of the following best describes 2-√4-x/4x as the limit approaches zero ?

OpenStudy (anonymous):

a. if fails to exist because it unbounded b. it fails to exist because its oscillates c it exist and equals to 1/16 d. it exist and equals to 2

OpenStudy (anonymous):

\[2-\sqrt{4-x/4x}\]

OpenStudy (anonymous):

or \[2-\sqrt{4-x}/4x\]

OpenStudy (anonymous):

the second one...

myininaya (myininaya):

are you sure it isn't this: \[\lim_{x \rightarrow 0}\frac{2-\sqrt{4-x}}{4x}\]?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(2-\frac{\sqrt{4-x}}{4x})\] if it really is this one... maybe we can combine the fractions and see what happens if we plug in 0

OpenStudy (anonymous):

\[\frac{ 2-\sqrt{4-x} }{ 4x } \times \frac{ 2+\sqrt{4-x} }{ 2+\sqrt{4-x} }\]

OpenStudy (anonymous):

sorry i was a way.... the first step is as @nelsonjedi wrote

OpenStudy (anonymous):

you don't need to multiply these together just the numerator and if look carefully you would know the answer.... the mid term should cancel out

OpenStudy (anonymous):

ido you still need help?

OpenStudy (jhannybean):

\[\lim_{x\rightarrow 0}(2-\frac{\sqrt{4-x}}{4x})\]\[\lim_{x\rightarrow 0} \left[\frac{2(4x)}{4x}-\frac{\sqrt{4-x}}{4x}\right]\]\[\lim_{x \rightarrow 0} \left[\frac{8x-\sqrt{4-x}}{4x}\right]\]\[\lim_{x\rightarrow 0} \left[ \frac{8x-\sqrt{4-x}}{4x} \cdot \frac{8x +\sqrt{4-x}}{8x +\sqrt{4-x}}\right]\]\[\lim_{x\rightarrow 0}\left[ \frac{16x^2-(4-x)}{4x(8x+\sqrt{4-x})}\right]\]\[\lim_{x \rightarrow 0} \left[\frac{16x^2 +x -4}{4x(8x+\sqrt{4-x})}\right]\]think you can simplify from here on out?